### v1

# Federated Automated Feature Engineering

[…]
main.bib
Federated Automated Feature Engineering
Tom Overman1 Diego Klabjan2
(1Department of Engineering Sciences and Applied Mathematics, Northwestern University
2Department of Industrial Engineering and Management Sciences, Northwestern University
1 Abstract
Automated feature engineering (AutoFE) is used to automatically create new features from original features to improve predictive performance without needing significant human intervention and domain expertise. Many algorithms exist for AutoFE, but very few approaches exist for the federated learning (FL) setting where data is gathered across many clients and is not shared between clients or a central server. We introduce AutoFE algorithms for the horizontal, vertical, and hybrid FL settings, which differ in how the data is gathered across clients. To the best of our knowledge, we are the first to develop AutoFE algorithms for the horizontal and hybrid FL cases, and we show that the downstream model performance of federated AutoFE is similar to the case where data is held centrally and AutoFE is performed centrally.
2 Introduction
Automated feature engineering (AutoFE) aims at automatically creating and selecting new features from the original set of features that improve downstream predictive performance. Feature engineering is often used to craft new features that help improve simple models. These crafted features may include nonlinear transformations of existing features that allow linear models to capture more complex, nonlinear interactions. Traditionally, feature engineering required specialized domain knowledge and extensive time for trial-and-error testing. However, recent approaches in automated feature engineering allow data scientists to automatically craft new informative features with very limited time, effort, and domain knowledge.
Federated learning (FL) is a privacy-preserving machine learning technique that has gained traction in recent years. The main requirement in federated learning is that data belonging to client devices cannot be directly sent to the server to build a global model. Instead, the typical approach is for each client to build a local model with only their own local data, then the clients can send their model weights to the server which aggregates these weights in various ways. Many of the key challenges in federated learning revolve around minimizing the number of communications between the clients and server and maintaining good performance when the data across clients are non-IID [fl].
There are three federated learning settings that correspond to how the data is gathered across clients. In horizontal federated learning, each client has a subset of samples and each of these samples has all of the features [fl]. In vertical federated learning, each client has the full set of samples but only a subset of the features [fedbcd]. Finally, in hybrid federated learning, each client only has a subset of samples and each of these samples only has a subset of features [hyfdca][hyfem][hybrid-graph]. Each of these settings requires different algorithmic approaches.
This paper focuses on automated feature engineering in the context of federated learning, where clients are not able to send data to the server, but we still build informative new features from the original raw features. We introduce algorithms for all three FL settings: horizontal, vertical, and hybrid. The horizontal algorithm allows each client to perform any AutoFE method of choice, the server aggregates the engineered feature strings, and then uses a communication-minimizing, hyperband-inspired selection method [hyperband] to determine which engineered features to keep. The hybrid and vertical FL methods hinge on homomorphic encryption to safely form combinations of features that reside on different clients, then use a similar selection scheme as the horizontal algorithm to determine the best engineered features to keep. All of the algorithms are iterative in that they gradually construct more and more complicated engineered features each iteration.
Our contributions are as follows
  1. 1.
    To the best of our knowledge, we are the first to develop federated AutoFE algorithms in the horizontal and hybrid FL settings. We also provide a new vertical AutoFE algorithm.
  2. 2.
    We provide experiments demonstrating that our horizontal FL AutoFE algorithm performs similarly to the case where AutoFE is performed centrally. This is a major achievement in FL where typically federated algorithms perform notably worse than the centralized counterparts.
In Section 3, we review the relevant work on both federated learning and AutoFE and discuss the minimal work that exist on AutoFE in the vertical FL setting. In Section 4, we detail our proposed AutoFE algorithms for the horizontal, vertical, and hybrid FL settings. In Section 5, we show results of our algorithm in the horizontal FL setting and demonstrate the performance is comparable to the case where data is located centrally and AutoFE is performed centrally.
3 Related Work
There is significant work existing on automated feature engineering. The main focus of most algorithms is on cleverly navigating the massive search space of possible combinations of features for high order engineered features. OpenFE uses a very fast boosting technique based on gradient boosted trees to quickly evaluate all possible feature combinations [openfe]. AutoFeat enumerates all possible combinations up to a specified order, then uses a fast feature selection technique to select the best performing features [autofeat]. IIFE tackles this problem by using interaction information to determine the best synergizing feature pairs and only exploring the best pairs of features, then iteratively builds higher order features [iife]. EAAFE uses a genetic algorithm approach to search the best engineered features [eaafe]. DIFER uses a deep learning approach to find representations of the engineered feature strings and their predicted validation scores [difer]. Furthermore, very recently there have been a focus on using Large Language Models to build engineered features in an informed fashion based on feature descriptions [caafe].
Federated learning also has extensive algorithmic work in addition to interesting theoretical results. Federated averaging is the original FL algorithm that is designed for the horizontal setting [fedavg]. Further improvements for heterogeneous data exist, such as SCAFFOLD [scaffold] and RadFed [radfed]. The vertical FL setting has also had significant work in recent years [fedbcd]. The hybrid federated learning case where clients have subsets of both samples and features has far fewer algorithms, but has recently had more work exploring this case [hyfdca][hyfem][hybrid-graph].
Despite the great deal of work in federated learning and automated feature engineering, there is relatively limited work on automated feature engineering in the federated learning setting. All of the existing work at the time of this writing focused on the vertical case, where each client has different features, and all clients have the same samples. There exists work in AutoML in a federated setting, however automated feature engineering is not one of the aspects discussed in this work [autofl]. FLFE is an algorithm for vertical federated AutoFE that uses quantile sketching to encode features and uses deep neural networks to predict if the transformation of features is useful [vertfefl]. The neural networks are trained on a large quantity of datasets. Finally, homomorphic encryption is used when constructing the best candidate features for building the final model. However, this work shows only small gains compared to using only the raw features, does not compare to AutoFE methods in the centralized setting, and does not provide any source code making reproducibility very difficult. Another recent work uses chi-square binning and homomorphic encryption to tackle the vertical federated AutoFE setting [vertfefl2]. However, this work does not share source code, only tests on two datasets, and only shares runtimes and not the actual test scores of using the new features. We believe this makes the work very difficult to reproduce and compare. It is clear that there is a large research gap in all three settings of federated AutoFE: horizontal, vertical, and hybrid. Our work provides algorithms for all of these settings.
4 Algorithms
The most straightforward approach to performing AutoFE in the federated learning setting is to simply run the AutoFE algorithm as usual but perform all model trainings and evaluations using federated algorithms. However, since most AutoFE algorithms require thousands of model trainings and evaluations and each of those model trainings would require many communications between clients and the server, the communication overhead makes this naive approach infeasible. The development of our federated AutoFE algorithms for each FL setting rely on minimizing this communication overhead.
4.1 Horizontal Setting
1 Input: Clients $k=1,2,...,Q$, Local data matrix $X_{k}\in\mathbb{R}^{N_{k}\times M}$ and targets $y_{k}\in\mathbb{R}^{N_{k}}$ belonging to client $k$, denote the list of feature strings of engineered features on client $k$ as $\mathcal{F}_{k}$, AutoFE routine $A_{k}:\{X_{k},y_{k}\}\rightarrow\mathcal{F}_{k}$ for each client $k$
2 Algorithm Parameters: Number of sparsity-refinement iterations $I$, Number of halving iterations $H$, number of starting candidates for $h$-th iteration $\mathcal{N}_{h}$, halving ratio $\gamma$, number of levels in $h$-th halving iteration $\mathcal{L}_{h}$, number of communications in level $l$ of $h$-th halving iteration $\mathcal{C}_{h,l}$, Number of outermost server iterations $J$, number of sparsities to select each iteration $s$
3 for each j=1,…,J do
4       for each client $k=1,2,...,Q$ do
5             $\mathcal{F}_{k}$ = $A_{k}(\{X_{k},y_{k}\})$
6             Send $\mathcal{F}_{k}$ to server
7      $\mathcal{F}$ = $\mathcal{F}_{1}\cup\mathcal{F}_{2}\cup...\cup\mathcal{F}_{Q}$
8       $s_{a}=0$, $s_{b}=1$
9       Server Executes:
10       for each sparsity-refinement iteration $i=1,...,I$ do
11             $\mathbb{L}=\{\;\}$
12             for each halving iteration $h=1,2,...,H$ do
13                   $\mathbb{L}$ = $\mathbb{L}\;\cup$ $\mathcal{F}$, $\mathcal{N}_{h}$, $\gamma$, $\mathcal{L}_{h}$, $\{\mathcal{C}_{h,l}\;\forall\;l\in 1,...,\mathcal{L}_{h}\}$, $s_{a}$, $s_{b}$)
14
15            Find $s$ best scores and corresponding sparsities of $g_{c}$ denoted as $\mathbb{S}_{top}$ in $\mathbb{L}$
16             Find mean $\mu_{\mathbb{S}}$ and standard deviation $\sigma_{\mathbb{S}}$ of sparsities in $\mathbb{S}_{top}$
17             $s_{a}=\max(0,\mu_{\mathbb{S}}-\sigma_{\mathbb{S}})$, $s_{b}=\min(1,\mu_{\mathbb{S}}+\sigma_{\mathbb{S}})$
18      Find best score and corresponding selection mask $g_{best}$ across all sparsity-refinement iterations
19       $\mathcal{F}=g_{best}\odot\mathcal{F}$
20       Each client builds new engineered numerical features $X_{k}$ from $\mathcal{F}$ and previous $X_{k}$.
Return $\mathcal{F}$
Algorithm 1 Horizontal-FLAFE
1 Input: List of feature strings $\mathcal{F}$, number of starting candidate masks $\mathcal{N}_{h}$, halving ratio $\gamma$, number of levels $\mathcal{L}$, number of communications in $l$-th level $\mathcal{C}_{l}$, sparsity lower bound $s_{a}$ and upper bound $s_{b}$
2 for each $c=1,2,...,\mathcal{N}_{h}$ do
3       Draw random sparsity $\mathcal{S}_{c}$ from Uniform($s_{a}$,$s_{b}$)
4       Generate random mask, $g_{c}\in\mathbb{R}^{|\mathcal{F}|}$, by drawing each element from a Bernoulli trial with probability of drawing 1 being $\mathcal{S}_{c}$.
[…]
6 for level $l=1,...,\mathcal{L}$ do
7       for each $g_{c}\in G$ do
8             Each client constructs the numerical features based on engineered feature strings $\mathcal{F}_{c}$ on local data
[…]
11
Return remaining $g_{c}\in G$ and their final scores $\text{score}_{\mathcal{L},g_{c}}$ as list of tuples $[(g_{c},\text{score}_{\mathcal{L},g_{c}})\;\forall\;g_{c}\in G]$
Algorithm 2 FedSuccessiveHalving
We first describe our algorithm for the horizontal FL setting; as far as we know, this is the first horizontal federated AutoFE algorithm to be developed. First, AutoFE is performed on each client separately only with local data on that client. This allows great flexibility in selecting an AutoFE algorithm, as the structure of Horizontal-FLAFE allows for any choice. Then, each client sends the feature string representation of the engineered features to the server. It is important to note that no data is actually sent to the server, only the string representation of the engineered feature such as log(X001**2/X002). The server then finds the union of all of the sets of engineered feature strings from each client. The union of the engineered feature strings is sent back to each client, where the clients then parse the feature string and compute the actual numerical feature vector.
The next step of the horizontal algorithm is selecting the best engineered features in a federated fashion. Our approach is inspired by Hyperband [hyperband] where successive halving is used with varying resource allocations (in this case the resource is communications between client and server). First, a large set of random candidate feature masks are generated. The feature masks are vectors of 0s and 1s that determine which features to keep and which to remove. The operation of masking the feature vector is denoted as $g\odot\mathcal{F}$. For each of these feature masks, each client removes the features specified by the feature mask and then the model is trained and evaluated using FedAvg (or any other horizontal FL algorithm) but with only a very limited number of communication rounds. Then, the top performing feature masks are kept, and the rest discarded. The next round repeats the same process but with a larger number of communication rounds in the FL training. This is repeated until the best performing masks are found with a large number of communication rounds.
This process is repeated for various number of levels similarly to Hyperband. Finally, we cannot simply uniformly draw our selection masks because this results in an average sparsity of 0.5. However, we find that depending on the dataset, base AutoFE algorithm used, and degree of non-IID across clients, the best performing sparsity on the feature mask greatly differs from 0.5. Thus, we have an outer loop that gradually refines the best performing sparsity. Finally, this entire process is repeated in an outermost loop in order to generate increasingly complex features as combinations of previously engineered features. However, the outputs from AutoFE are generally already fairly complex, so the number of outermost iterations is usually kept to a low number.
Figure 1: Workflow of one outer iteration of Horizontal-FLAFE with 3 clients and just a few engineered features. Typically there are more clients and each local AutoFE process generates many more complicated engineered features.
4.2 Vertical Setting
The basic idea for the vertical algorithm is that for features that are on separate clients and can’t be directly shared, we need to homomorphically encrypt the vectors, combine the features in the encryped space, encrypt and then add Laplace noise, and finally send the combined features back to the clients to decrpyt. This combined feature can then be tested to see if it is a useful new feature.
The only homormorphic encryption schemes that are fast enough for practical use preserve addition and multiplication operations. Thus, any complicated functions must be decomposed into a Taylor series of additions and multiplications. However, since homomorphic encryption is very costly, especially for large datasets, we want to minimize the number of times we need this operation. This is why we first find the feature importance of each feature held on the client, and only the most important features on this client are combined with the most important features on other clients.
Once all of the engineered features are created and sent back to clients, the best engineered features are selected in the same way that the feature selection mask is found in Algorithm 1. The successive halving steps require training and evaluating in the vertical FL setting, so vertical FL training algorithms are needed. It is important to note that most homomorphic encryption schemes require specifying the downstream function while encrypting the inputs, while in our approach we apply various functions (combinations of multiplications and additions) to the encrypted inputs. Fortunately, BFV homomorphic encryption does not require the specification of the downstream function while encrypting and fits well within our approach [bfv1][bfv2]. This process is shown in Figure 2.
1 Input: Clients $k=1,2,...,Q$, Local data matrix $X_{k}\in\mathbb{R}^{N\times M_{k}}$ and targets $y_{k}\in\mathbb{R}^{N}$ belonging to client $k$, BFV Homomorphic encryption function $H(\cdot)$
2 $\mathcal{F}_{final}=[\;]$
3 for each outer iteration $i=1,...,O$ do
4       for each client $k=1,...,Q$ do
5             $\mathcal{F}_{k}=[\;]$
[…]
6             Find $K$ most important features on client $k$ denoted as $\mathcal{F}_{important,k}\subseteq\mathcal{F}_{k}$
7             $\mathcal{H}_{k}=[\;]$
8             for each $f\in\mathcal{F}_{important,k}$ do
9                   Homomorphic encrypt feature vector denote as $H(f)$
10                   $\mathcal{H}_{k}=\mathcal{H}_{k}\cup H(f)$
11            Send list of encrypted feature vectors on client $k$, $\mathcal{H}_{k}$, to server
12      Server executes CreateAndSendNonLocalCombinations()
13       Clients receive new combined features from server
14       Each client decrypts the features they were sent $F_{i,j}=\text{Dec}(H_{i,j})$ back to the original feature space
15       Select best newly engineered features, denoted as $\mathcal{F}_{new}$, using Lines 6-14 of Algorithm 1, this requires Vertical FL training algorithms
16       $\mathcal{F}_{final}=\mathcal{F}_{final}\cup\mathcal{F}_{new}$
Return $\mathcal{F}_{final}$
Algorithm 3 Vertical-FLAFE
1 Input: features F where $i$-th column is indexed as $F_{i}$, set of candidate bivariate transformations
2 $F_{out}=[\;]$
3 for each $(F_{i},F_{j})\in F\times F$ do
4       for each candidate bivariateFunction() do
5             $F_{out}=F_{out}\cup\text{bivariateFunction}(F_{i},F_{j})$
6
Return $F_{out}$
Algorithm 4 CreateLocalCombinations()
1 Input: Lists of encrypted feature vectors for each client denoted as $\mathcal{H}_{k}$, set of candidate bivariate transformations, $b$ is the scale parameter used for the Laplace noise
2 $\mathcal{H}=\mathcal{H}_{1}\cup...\cup\mathcal{H}_{Q}$
3 $H_{out}=[\;]$
4 for each $(H_{i},H_{j})\in\mathcal{H}\times\mathcal{H}$ do
5       for each candidate bivariateFunction() do
             $H_{i,j}=\text{bivariateFunction}(H_{i},H_{j})+H(\text{Lap}(0|b)^{N})$ // inject Laplace noise
6             Select client that owns feature $i$ or feature $j$ at random with equal probability of selecting between the two
7             Send $H_{i,j}$ back to selected client
8
Algorithm 5 CreateAndSendNonLocalCombinations()
Figure 2: Workflow of Vertical-FLAFE with 3 clients and each have two starting features. Typically there are more clients and the allowed function transformations are more complex.
4.3 Hybrid Setting
We make some assumptions on the way that data is gathered across clients in the hybrid setting. While work such as [hyfdca] allows fairly flexible partitions of data as long as each client hold rectangular patches of the overall data, we enforce more strict rules on the way data is split. Specifically, we require that every sample be partitioned across clients in the same way. Essentially, the vertical divisions must run across the entire sample space without any jagged partitions. An example of a suitable partition is shown in Figure 3. Algorithm 6 follows very similarly to Algorithm 6, but has some key differences that are needed in the hybrid setting. First, the process of finding the most important features on each client to combine with other client’s features needs to be done across several clients because each feature is held across many clients in the hybrid setting. This likely means training with horizontal FL across clients that share the same features but different samples to determine feature importances. Furthermore, selection of the best generated features requires hybrid FL training algorithms, of which there are very few and these typically have assumptions on the type of models they work for. For example, as far as we know, there are no hybrid FL algorithms for tree-based models.
Figure 3: Example of how data is assumed to be partitioned for Hybrid-FLAFE.
1 Input: Clients $k=1,2,...,Q$, Local data matrix $X_{k}\in\mathbb{R}^{N_{k}\times M_{k}}$ and targets $y_{k}\in\mathbb{R}^{N_{k}}$ belonging to client $k$, BFV Homomorphic encryption function $H(\cdot)$
2 $\mathcal{F}_{final}=[\;]$
3 for each outer iteration $i=1,...,O$ do
4       for each client $k=1,...,Q$ do
5             $\mathcal{F}_{k}=[\;]$
[…]
6             Find $K$ most important features on client $k$ denoted as $\mathcal{F}_{important,k}\subseteq\mathcal{F}_{k}$, perform this across all clients that contain the specified feature (may require FL training depending on chosen feature selection method)
7             $\mathcal{H}_{k}=[\;]$
8             for each $f\in\mathcal{F}_{important,k}$ do
9                   Homomorphic encrypt feature vector denote as $H(f)$
10                   $\mathcal{H}_{k}=\mathcal{H}_{k}\cup H(f)$
11            Send list of encrypted feature vectors on client $k$, $\mathcal{H}_{k}$, to server
12      Server executes CreateAndSendNonLocalCombinations()
13       Clients receive new combined features from server
14       Each client decrypts the features they were sent $F_{i,j}=\text{Dec}(H_{i,j})$ back to the original feature space
15       Select best newly engineered features, denoted as $\mathcal{F}_{new}$, using Lines 6-14 of Algorithm 1, this requires Hybrid FL training algorithms
16       $\mathcal{F}_{final}=\mathcal{F}_{final}\cup\mathcal{F}_{new}$
Return $\mathcal{F}_{final}$
Algorithm 6 Hybrid-FLAFE
5 Experiments
We conduct basic experiments to evaluate how well Horizontal-FLAFE performs compared to standard AutoFE in the centralized case. We perform experiments on the OpenML586 and Airfoil datasets, the base AutoFE algorithm used is AutoFeat with 2 feature engineering steps. The downstream model used is Lasso with a regularization constant of 0.01. We set $J=1$, $I=1$, $H=1$, and $\mathcal{C}_{1,1}=16$. We split the data across 8 clients in an IID fashion. The clients hold 60% of the overall dataset, and the server holds 20% as a final hold-out test set, and 20% as a validation set used in determining the best feature mask. The models are trained in a federated fashion using FedAvg.
We compute the scores for AutoFE with fully centralized data and observe how close the FL approaches get to this score. Finally, we also compute the test scores for the case with no feature selection mask (the full union of engineered features is used), in order to justify the process of selecting the best feature selection mask.
Figure 4 shows the results on these two datasets. The plot shows the baseline test score, where the model is trained centrally and only the original, raw features are used. The centralized test score is such that AutoFE is performed centrally, and then, the final model is trained centrally with the best engineered features. The “without mask” bar references the final model trained with FedAvg, where all of the engineered features found on each client are used without performing any selection. Finally, the Horizontal-FLAFE bar shows the test score with the final model trained with FedAvg and the best engineered features created and selected using Algorithm 1. It is important to note that we do not compare with any other Horizontal federated AutoFE approaches, because as far as we know, no other horizontal federated AutoFE algorithms exist.
These results demonstrate that Horizontal-FLAFE approaches the same test scores as if AutoFE is performed centrally, and then, the final model is trained centrally. In fact, the test score for Airfoil is higher for Horizontal-FLAFE than centralized AutoFE. We believe this is due to the fact that Horizontal-FLAFE creates many more engineered features as the union is taken across all clients, and we do not expect this to hold for non-IID experiments. Furthermore, there is a clear increase in test score between the optimal selection mask and without a selection mask (using the full union of engineered features); this justifies our resource-aware selection mask optimization. The best feature mask for OpenML586 had a sparsity of 0.902 and for Airfoil the best sparsity was 0.708, indicating that our sparsity refinement approach is better than just uniformly randomly selecting the candidate feature masks.
Figure 4: Performance of Horizontal-FLAFE on two different datasets.
6 Conclusions and Future Work
We contribute three AutoFE algorithms that operate in the three key federated learning settings: horizontal, vertical, and hybrid. We provide experimental evidence that Horizontal-FLAFE performs similarly to performing the AutoFE process in a centralized manner. Our future work is focused on expanding the experimental results across more datasets and choices of algorithm hyperparameters and having extensive experiments performed in the vertical and hybrid FL settings.
[…]